how many ounces of pure water must be added to 60 ounces of a solution that is 20% water to make a solution that is 30% water
Answer by mananth(12270)

let x ounces = water
x+60*0.2=(x+60)*0.3
x+12=0.3x+18
0.7x=6
x=6/0.7
x=8.57
...m.ananth@hotmail.ca